### v1

# Unintended Impacts of LLM Alignment on Global Representation

Abstract
Before being deployed for user-facing applications, developers align Large Language Models (LLMs) to user preferences through a variety of procedures, such as Reinforcement Learning From Human Feedback (RLHF) and Direct Preference Optimization (DPO). Current evaluations of these procedures focus on benchmarks of instruction following, reasoning, and truthfulness. However, human preferences are not universal, and aligning to specific preference sets may have unintended effects. We explore how alignment impacts performance along three axes of global representation: English dialects, multilingualism, and opinions from and about countries worldwide. Our results show that current alignment procedures create disparities between English dialects and global opinions. We find alignment improves capabilities in several languages. We conclude by discussing design decisions that led to these unintended impacts and recommendations for more equitable preference tuning.
Unintended Impacts of LLM Alignment on Global Representation
Michael J. Ryan Stanford University michaeljryan@stanford.edu William Held Georgia Institute of Technology wheld3@gatech.edu Diyi Yang Stanford University diyiy@cs.stanford.edu
1 Introduction
Recently, LLM-powered chat assistants OpenAI (2023a); Touvron et al. (2023); Tunstall et al. (2023b) have exploded in popularity. As of December 2023, ChatGPT has amassed over 100M weekly users OpenAI (2023b) and Llama-Chat-7B is downloaded almost one million times a month from HuggingFace111Llama-Chat-7B Huggingface Page. The success of these chat models is dependent on "alignment", which takes a base model with a language modeling objective and produces an instruction following preference-guided model to better serve user interests. Practitioners use algorithms such as RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2023) to optimize models for attributes such as helpfulness and harmlessness and give them their chat assistant persona Ouyang et al. (2022); Bai et al. (2022); Zhu et al. (2023).
Unlike the nebulous pre-training process, which is largely defined by the distribution of data online Raffel et al. (2019); Gao et al. (2020); Computer (2023), model developers have a high degree of control for the key alignment variables. Who will give feedback? What prompts/tasks are in-domain? Who will provide exemplar responses? These are just a few design decisions that underscore a larger question: Whose preferences are we aligning LLMs with, and crucially, whose preferences are we missing? As Blodgett et al. (2020) put it, "For which speakers are NLP systems developed?"
This question often does not have an explicit answer in current alignment practices Bakker et al. (2022), making it unclear which model behaviors are intentional normative judgments and which are unintended biases. For example, the Starling 7B Reward Model Zhu et al. (2023) gives higher scores to responses claiming to be from English-speaking Western nations and lower scores for Middle Eastern and African nations (See Figure 1). In this work, we take a closer look at the effects these design decisions have on a model’s ability to serve a global population, which is key to understanding if the general use of aligned LLMs Eloundou et al. (2023) is likely to be positively adopted globally.
Existing performance evaluations of chat assistants mainly focus on tasks such as reasoning Clark et al. (2018); Zellers et al. (2019); Sakaguchi et al. (2021); Cobbe et al. (2021), multitask knowledge Hendrycks et al. (2021); Suzgun et al. (2023), truthfulness Lin et al. (2022), multi-turn instruction following Zheng et al. (2023), and similar variations of broad knowledge/reasoning/skills Chen et al. (2021); Zhong et al. (2023). Instead, we explore a set of representative domains covering variations common in diverse global user bases: English dialects, multilingualism, and global opinions, and show a direct impact on model performance.
Our evaluations focus on measuring how alignment makes LLMs more agreeable and helpful for different groups of possible global users. While prior works (§2) have explored the representation of global opinions in language models (Durmus et al., 2023; Santurkar et al., 2023), they only study the final model. However, the process of transforming a base language model to a user-facing chat model involves two key sequential steps: supervised fine-tuning (SFT) and preference tuning (PT). The impacts of alignment are the product of the base model, SFT, and PT. In addition to evaluating surveys, we study performance gaps on downstream tasks that occur throughout the alignment process for several variations common in global user bases. Together, these evaluations assess whether alignment procedures make LLMs both more agreeable and helpful for a global user base. In summary, our contributions are as follows:
1.
We first evaluate the effects of alignment in a purely English setting, focused on global dialects of English (§4). Effective alignment procedures improve performance on an English intent prediction task for conversations between US, Indian, and Nigerian speakers (Eisenstein et al., 2023). However, alignment significantly increases the disparity between English dialects from about 1% before alignment to as high as 17.1% after alignment.
2.
We then evaluate the effects of alignment on model multilingualism (§5). Despite most models branding themselves as primarily English, alignment largely improves multilingual performance in two question-answering tasks, highlighting a positive unintended impact.
3.
Finally, we evaluate the effects of alignment on a model’s correlation with global opinions from particular countries and about particular countries (§6). We find that all evaluated alignment procedures increase the similarity between model responses and opinions from the US relative to major nations from other regions, such as China, Jordan, and Nigeria. We further release a new dataset of 554 opinionated questions about countries from r/AskReddit. We find that the open-source Starling reward model, on average, rates 99.4% of all other countries more negatively than the USA. However, this bias does not seem to propagate to the language model preference-tuned with this reward model.
2 Related Work
Large Language Model Biases.
Several works have explored various biases in large language models Ferrara (2023). Specifically, prior work has explored dialect bias Ziems et al. (2023), language bias Nicholas and Bhatia (2023); Yong et al. (2023), political bias Santurkar et al. (2023); Hartmann et al. (2023), cultural bias Naous et al. (2023); Durmus et al. (2023); Huang and Yang (2023), gender bias Kotek et al. (2023); Treude and Hata (2023); Wan et al. (2023), and more Nadeem et al. (2021); Cao et al. (2023); Dhingra et al. (2023). Though some LLMs studied in these works underwent RLHF or SFT, these works do not directly measure the bias introduced by the alignment process. In contrast, our study seeks to identify the unintended impacts exacerbated by alignment.
Negative Impacts of Preference Tuning.
Lambert et al. (2023) provides a fantastic overview of the risks of RLHF. Prior work has noted the social impacts of RLHF and Preference Tuning Liu (2023). Ouyang et al. (2022) identifies that they aligned their model with mostly English speakers from the US and Southeast Asia. RLHF has been observed to steer models towards outputs that are longer Singhal et al. (2023), more assertive Hosking et al. (2023), and less novel Kirk et al. (2023). RLHF can also make mistakes in the output more subtle Bai et al. (2022). Shaikh et al. (2023) finds that RLHF decreases grounding acts. Perez et al. (2023) find that RLHF makes models echo user opinions, stronger political views, and requests not to be shut down.
Santurkar et al. (2023) perform the exploration most similar to our work. The authors investigate how base and post-RLHF models differ in political opinions with 60 USA demographic groups. Our study expands this experimentation beyond surveying LLMs to assessing downstream performance on various tasks. We also investigate global opinions and values outside US demographics.
| Model | Preference-Tuning | Feedback | Preference Data | SFT Model | SFT Data | Base Model | Pre-training Data |
| Llama 2 Chat | PPO (RLHF) | Human | Proprietary | – | Proprietary | Llama 2 | Internet Dump* |
| Tulu 2 DPO | DPO | GPT-4 | UltraFeedback | Tulu 2 | Mixed | Llama 2 | Internet Dump* |
| Starling LM | PPO (RLAIF) | GPT-4 | Nectar | OpenChat 3.5 | Mixed | Mistral v0.1 | Internet Dump* |
| Zephyr Beta | DPO | GPT-4 | UltraFeedback | Mistral SFT | UltraChat | Mistral v0.1 | Internet Dump* |
3 Alignment Process
First, we identify models with checkpoints at different stages of the alignment process (See Figure 2) so that we can measure the effects of each stage.
Supervised Fine-tuning.
In the supervised fine-tuning stage, the model is provided with prompts and example completions and fine-tuned to produce these sorts of completions. Popular SFT datasets for chat models include the human-written Flan222Note that Flan contains templated completions of other datasets rather than being fully naturally written Wei et al. and Open Assistant Köpf et al. (2023) datasets, and the synthetic ShareGPT 333https://sharegpt.com, Alpaca Taori et al. (2023), and Open-Orca Lian et al. (2023) datasets. All are variants of instruction following completions to task-oriented prompts. Typically, this step is used to make language models follow instructions rather than continue the input text based on the language modeling objective.
Preference Tuning.
After SFT, models undergo preference tuning, where a dataset of prompts and preference-ranked completions are used to align LLMs with user preferences. Two popular algorithms for preference tuning are Proximal Policy Optimization (PPO) Schulman et al. (2017), which is used in Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022), and Direct Preference Optimization (DPO) Rafailov et al. (2023). For RLHF, a reward model is trained, which takes in a prompt and completion and outputs a score predicting the degree of human preference for such an output, whereas, in DPO, the model is updated directly using the preference dataset.
Deployment
After alignment, language models are either deployed inside a product or released for broader use. Notably, these models are a core technology that enables higher-level user-facing systems. While model developers may intend a specific audience, open-access models can be adopted anywhere and major LLM APIs are globally accessible444OpenAI and Google Supported Countries. As a result, due to the broad nature of their possible utility, even unintended impacts of alignment can affect their global adoption.
Model Selection
We experiment on 9 distinct LLMs from two main model families licensed for academic use: Llama 2 7B Touvron et al. (2023) and Mistral v0.1 7B Jiang et al. (2023). We specifically focus on four distinct chat models built on these base models: Llama 2 7B Chat Touvron et al. (2023), Tülu 2 7B DPO Ivison et al. (2023), Starling LM 7B Zhu et al. (2023), and Zephyr-7B-beta Tunstall et al. (2023b). Each model underwent both SFT and preference-tuning. We explore all intermediate SFT models except for Llama 2 Chat since the SFT model has not been released. The SFT models for Tülu 2 7B DPO, Starling LM 7B, and Zephyr-7B-beta are Tülu 2 Ivison et al. (2023), OpenChat3.5 Wang et al. (2023), and Mistral-7B-SFT-beta Tunstall et al. (2023a) respectively. These models cover a variety of preference-tuning algorithms, feedback sources, and datasets. An overview of the models included in our study can be found in Table 1. We include prompts and other model details in Appendix B.
4 Global Representation: English Dialects
We first explore how preference tuning affects global English dialects by looking at model performance on a dialogue intent prediction task for three groups of global English speakers: US American, Nigerian, and Indian.
Task Setting
We experiment using the Multi-dialect Dataset of Dialogues (MD3) Eisenstein et al. (2023). MD3 is a high-quality collection of task-oriented transcripts from global English speakers. For MD3, we explore the intent prediction task for American English, Indian English, and Nigerian English speakers. In MD3, one player gives hints to the other to help them guess a secret word or “intent” without using any of the “distractor” words. To restrict to achievable inputs, we filter out any transcripts where the participants report failing to guess the correct intent. The language model is used to predict the intent of the dialogue using a brief description of the game and the transcript truncated right before the correct guess. We take a successful language model guess to be the case where the correct answer is generated, and no distractor words are generated.
Alignment Improves Performance in all Dialects, but Increases Disparity Between Dialects
We report the different accuracies of LM guesses in Figure 3 with 95% confidence intervals. Whenever changes to performance are significant (p<0.05), the alignment steps increase US English performance much more significantly than other global Englishes. Before alignment, all Base models performed relatively the same across dialects (about 5% accuracy for Llama and 8% accuracy for Mistral). Though SFT improves performance across all dialects it creates a disparity in performance gains between dialects. For Mistral SFT, the difference between USA English and Indian English grows from 0.98% in Mistral to 15.0% in Mistral SFT, and for USA English and Nigerian, the difference grows from 1.3% to 10.3%. Similarly, for OpenChat, the disparity in USA English and Indian English grows from 0.98% to 16.9% and between USA and Nigerian English from 1.3% to 9.7%.
Changes due to PT are far less impactful. However, in the case of Mistral SFT to Zephyr, the USA change is significantly positive. For OpenChat to Starling, the changes are not significant, but it is worth noting that the decrease in the correct answer rate in Nigeria is the largest. This suggests that PT also improves the disparity between US English and other dialects.
5 Global Representation: Languages
We investigate global language representation by measuring the multilingual ability of aligned LLMs on extractive QA and reading comprehension. We explore nine typologically diverse languages.
Task Setting
We utilize the Typologically Diverse Question Answering (TyDiQA) dataset Clark et al. (2020) to assess the multilingual capabilities of the LLMs. Specifically, we use the TyDiQA Gold Passage (GoldP) task, a collection of questions and single-paragraph passages spanning nine typologically diverse languages: Arabic, Bengali, English, Finnish, Indonesian, Korean, Russian, Swahili, and Telugu. The goal of the GoldP task is to extract the correct answer span from the passage. We assess models in the 1-shot setting by randomly sampling a demonstration from the train set, and we use greedy decoding for answer generation. We assess generated answers using CFM scores Li et al. (2024), a trained classifier over F1 scores and similar text features, which has been shown to correlate well with expert judgments.
To measure multilingual understanding, we use the Belebele benchmark Bandarkar et al. (2023), a parallel dataset of reading comprehension multiple-choice questions in 122 language variants. The dataset includes 900 questions per language variant written about 422 distinct passages from the Flores-200 Team et al. (2022) parallel dataset. We filter to the nine TyDiQA languages for comparison. We use language modeling probability on letter choices (A) to (D) to assess the model selection.
Alignment for English can improve Multilingual Performance.
Despite the stated goal to create English chat assistants, we find gains across most languages after alignment. For the reading comprehension task, we observe significant improvements across most languages and never a significant decrease in performance. For the TyDiQA extractive QA task, both Tülu and Starling improved in most languages. Zephyr TyDiQA performance decreases significantly in six of nine languages. All models worsen in Bengali to varying degrees: 12.7% worse for Llama Chat, 8.2% worse for Tülu, 9.7% worse for Zephyr, and 0.8% worse for Starling.
| Language | Tülu SFT | (%) | UltraChat | (%) |
| English | 1,146,844 | 86.9 | 1,458,969 | 99.9 |
| Spanish | 33,091 | 2.5 | 876 | 6.0E-4 |
| French | 30,977 | 2.3 | 359 | 2.5E-4 |
| Korean | 23,293 | 1.8 | 4 | 2.7E-6 |
| Japanese | 20,926 | 1.6 | 9 | 6.2E-6 |
| German | 12,270 | 0.93 | 65 | 4.5E-5 |
| Portuguese | 9,376 | 0.71 | 23 | 1.6E-5 |
| Russian | 9,137 | 0.69 | 13 | 8.9E-6 |
| Italian | 7,342 | 0.56 | 33 | 2.3E-5 |
| Indonesian | 3,761 | 0.29 | 3 | 2.0E-6 |
Multilinguality in Tülu SFT data Explains the Improvement in Multilingual QA Performance.
We run language identification to detect the languages that comprise the OpenChat and Tülu SFT datasets. Details on the language ID systems used are provided in Appendix D. Language ID results for the Tülu SFT data mix and UltraChat dataset for Zephyr are reported in Table 2. Although the full SFT split of OpenChat was not released, the authors also mention training on ShareGPT, Open Orca, and Open Assistant, so it overlaps with the Tülu SFT data mix through those sources.
Despite the intentions of Ivison et al. (2023) to train Tülu on English data, the Tülu SFT data is quite multilingual. In fact about 13.1% of the dataset is non-English. This explains the impressive improvement of the Tülu SFT model on Belebele and TyDiQA for most languages. Language ID also explains the decrease in Bengali performance. We find just 71 examples of Bengali in the Tülu SFT data (comprising of the data) and 0 examples of Bengali in UltraChat. Tracing the source of the multilingual data the Tülu data mix we find 141,970 non-English samples from ShareGPT, 16,801 samples from FlanV2, and 11,441 samples from Open Assistant.
The OpenChat Model (SFT between Mistral and Starling), like Tülu, also has impressive Multilingual gains, likely due to the overlapping use of ShareGPT and Open Assistant. UltraChat, on the other hand, seems to have gone through a more aggressive filter, which limits 99.9% English. While Llama Chat does not detail the SFT data, the explicit English focus of the model development makes it likely that the proprietary dataset is similarly curated. This explains the decrease in multilingual performance for Mistral SFT and Llama Chat in most languages for TyDiQA.
6 Global Representation: Opinions
The final axis of global representation we measure is global opinions. We measure LLM agreement with countries’ opinions on polarizing questions.
Task Setting
For measuring alignment with global values, we use GlobalOpinionsQA Durmus et al. (2023), a dataset of 2,556 questions and answers from cross-national surveys on global issues. The dataset contains distributions of responses from representative samples of over 100 nations with topics such as politics, media, technology, religion, race, and ethnicity. However, most questions in GlobalOpinionsQA are asked to only a few countries. To evaluate relative alignment between regions, we take the countries with the most responses from Asia, Europe, the Middle East, North America, South America, Oceania, and Sub-Saharan Africa. We then filter to questions with responses from all seven countries. This results in 245 questions, with answers from representative samples of The USA, China, Jordan, Brazil, Nigeria, Germany, and Australia.
We measure the probability of responding with each answer choice and compare the probability distribution with global respondents. Following the GlobalOpinionsQA paper Durmus et al. (2023) we measure 1-Jensen Shannon divergence between the LLM responses and responses for each country. We use a similar task setting to the original Anthropic paper. However, our analysis covers nine open models and all alignment stages, while the original analysis is limited to the Claude model. We report the change in similarity to global values in Figure 5 with 95% confidence intervals.
Alignment increases relative agreement with the USA versus Jordan (MENA), China (Asia), and Nigeria (SSA).
Our findings on GlobalOpinionsQA showcase that aligned language models tend to agree more closely with USA opinions than base language models. From Llama to Llama Chat, the difference between the USA similarity increases from 0.3% to 4.5% for Jordan, from 1.4% to 3.1% for China, and from -2.5% to 3.5% for Nigeria, showing around a 2-5% relative decrease in agreement versus the United States. For Western Nations like Germany or Australia, however, the agreement does not significantly change with respect to the USA. Similar trends hold for all the models. Interestingly, all models go from agreeing more with Nigeria than the USA before alignment to agreeing more with the USA than Nigeria after alignment. Our results agree with the findings of Durmus et al. (2023) that LLMs align to Western preferences, and show that this is exacerbated by alignment.
6.1 Reward Model Probing
GlobalOpinionsQA provides a rich testbed for measuring LLM agreement to opinions of certain countries, but it does not enable exploring opinions about specific countries. To better understand these learned opinions about countries, we explore the preferences of an Open Source Reward Model. We probe the Starling 7B Reward Model Zhu et al. (2023) and explore how its preferences vary on several counterfactual country opinion-based questions rather than multiple-choice questions.
Data Collection
Reward models are not well suited for multiple-choice assessments due to the limited response length. We build a dataset suited to counterfactual reward probing by collecting a set of 554 country-specific questions from the subreddit r/AskReddit 555https://www.reddit.com/r/AskReddit/. We search for questions using the queries "Which Country", "What Country", "Best Country", and "Worst Country" to collect varied questions. This resulted in 957 questions. After removing duplicates, questions with strictly factual answers, and questions that could not be answered with a specific country name, we were left with 554 quality-assured questions.
Two authors manually labeled each question as "positive" or "negative," where the positively labeled examples reflected something good about a country, and the negatively labeled indicated something bad. For example, "Which country do you never want to visit?" has a negative label, and "Which country has the best flag?" has a positive label. After independent labeling, the authors had a Cohen’s kappa of 0.963, disagreeing on just 10 labels, which were resolved after discussion.
We use ChatGPT to write completion templates for each question and manually validate their quality. For instance, the question "Which country has the best flag?" has the response template "{country} has the best flag, in my opinion." Finally, we categorize the questions into 11 categories: "Aesthetics," "Cuisine," "Culture," "Geopolitics," "History," "Personal," "Preferences," "Quality of Life," "Speculation," "Stereotypes," and "Tourism." More details and examples can be found in Appendix C.
Task Setting
We probe the Starling 7B Reward Model with all 554 questions and 181 countries with a population over 250,000 to fill in as answers. For each question we record the score assigned by the reward model to each country. Since reward models are primarily used for pairwise comparisons, we use the RM to assign a rank to each country per question based on the outputted reward. We then compute the mean rank for each country over all questions. We invert the rankings on "negative" questions, so a low ranking is always preferable.
| Country | Starling RM | US Citizens | ||
| Rank | Final | Mean | 2017 | 2023 |
| UK | 1 | 67.6 | 2 | 2 |
| Canada | 2 | 76.1 | 1 | 1 |
| Japan | 3 | 77.2 | 3 | 4 |
| France | 4 | 78.1 | 4 | 3 |
| India | 5 | 84.4 | 6 | 7 |
| Palestine | 15 | 111.9 | 14 | 13 |
| Russia | 16 | 113.9 | 13 | 18 |
| Iraq | 17 | 120.0 | 16 | 14 |
| Afghanistan | 18 | 129.1 | 17 | 15 |
| North Korea | 19 | 152.1 | 19 | 19 |
Starling RM Correlates with US opinions
We measure correlation with rankings by US citizens collected from Gallup polls in 2017 and 2023 Brenan (2023). Gallup surveyed 1,035 US adults in 2017 and 1,008 US adults in 2023 and asked them to rate countries as "Very Favorable," "Mostly Favorable," "Mostly Unfavorable," "Very Unfavorable," or "No opinion." The aggregate scores are used to compute a ranking over the 21 countries surveyed. We report the top 5 and bottom 5 countries from this list ranked by Starling in Table 3. Comparing just the rankings of these 21 countries to those produced by the Starling 7B RM, we find a 0.926 Spearman correlation with the 2017 results and a 0.849 Spearman correlation with the 2023 results. This indicates a high overlap between US opinions and the learned preferences of the Starling RM. These results offer a step towards answering the question, "To whose preferences are we aligning language models?" Western preferences certainly have a significant influence. We report all rankings by Starling RM along with a choropleth visualization in Appendix E. Unrestricted to the Gallup list, Starling ranks "Morocco," "the USA," "Slovenia," and "New Zealand" highest and "Western Sahara," "North Korea," "Turkmenistan," and "Central African Republic" lowest. Similar to our motivating example, "Where are you from?" we find the Starling model assigns low rewards to countries in central Africa and the Middle East.
Reward Models have Little Influence on Out-of-Distribution Preferences
We compute rankings of all countries using perplexity on the same questions for all models. We report Spearman rank correlation in Figure 6. Within the model families (Llama vs Mistral), rankings vary only slightly. Llama, Llama Chat, Tulu SFT, and Tulu DPO correlate highly, and Mistral, Mistral SFT, Zephyr, OpenChat, and Starling LM all correlate tightly. Interestingly, Starling RM predictions correlate poorly with all models, including Starling LM, suggesting the preferences were not reflected in the model. This case study raises a fascinating finding: the pre-training data defines the model behavior on out-of-distribution preferences. If opinionated country questions don’t show up in the preference-tuning process, the reward signal does not steer the LLM, and it retains the preferences of the base model.
7 Discussion and Conclusion
Our findings underscore three key recommendations for practitioners aligning LLMs.
The Alignment of Language Models is not a One-Size-Fits-All Solution.
Various groups are impacted differently by the alignment procedure. Transparency is of the utmost importance in disclosing the design decisions that go into aligning an LLM. Each step of alignment adds additional complexities and impacts on end users. As such, transparent reporting Mitchell et al. (2019); Bommasani et al. (2023); Longpre et al. (2023); Liesenfeld et al. (2023); Gilbert et al. (2023) ideally should encompass the entire alignment pipeline, not just the final model. The InstructGPT paper Ouyang et al. (2022) reports the demographics of their preference annotators, but most human-written preference datasets since then have not. Reporting such information, along with decisions about what prompts or tasks are in the domain, is essential for the responsible dissemination of aligned LLMs to a diverse audience of users Sorensen et al. (2024).
Slightly Multilingual SFT Data can have an Outsized Impact.
We find that just 13.1% of the Tülu dataset is in any language other than English, and yet this multilingual data leads to performance improvements in six out of nine tested languages for extractive QA and all nine languages for reading comprehension. On the reading comprehension task, we still see the greatest gains in English for Tülu, indicating this is not a trade-off but that many languages can benefit from multilingual data.
Reward Models do not Shape Model Preferences on Out-of-Distribution Settings.
When probing the Starling RM, we find a high correlation to the USA’s opinions of other countries. However, when we explore whether the models share these preferences, we find little correlation between the two. The similarity in country preferences is instead mostly consistent between model families. This suggests that for out-of-distribution settings such as this country-opinion domain, reward models do not influence the model they are tuning. This highlights that, beyond the reward model itself, the selection of the original SFT data and of PT prompts significantly shape the possible impacts of PT.
In conclusion, we identified three axes of global representation that are impacted by the alignment of language models: English dialects, multilingualism, and global opinions. From the mixture of training data to annotator demographics, many decisions go into aligning language models. We shed light on how some of these decisions can unintentionally impact global representation.
Limitations
In this paper, we explore nine open-source language models at various alignment stages on four downstream tasks. Since Llama 2 SFT has not been publicly released, we cannot disentangle the effects of SFT and RLHF in the alignment of Llama 2 Chat. We use the released model checkpoints on Huggingface for all of the open-source models tested in this paper. Since we use open checkpoints rather than aligning the models ourselves, we cannot directly test individual changes to the alignment procedure and their downstream impacts. Instead, we focus a wider lens on the practical downstream effects of each alignment stage. We leave causal intervention and interpretability studies on the impacts of alignment to future work.
We select our datasets based on high-quality natural human-written benchmarks. Based on the availability of such high-quality resources, we focus on intent detection for dialects, extractive QA and reading comprehension for languages, and global opinion surveys for opinions. Since we test on a limited set of tasks, it is possible that failure modes arise on tasks that we did not assess in this work. In the context of our multilingualism experiments, we find that the performance improvements in all languages span two tasks. A more concrete assessment of multilingual generalization would benefit from a wider breadth of tasks.
Ethics Statement
In our discussion of LLM multilingualism and dialect support, we make the normative assumption that it is positive for LLM to express greater capabilities in these languages and language varieties. This operates under the assumption that the subsequent deployment of said technologies in the real world will be a process that is done with and for speakers. However, we acknowledge that this is frequently untrue and that technology such as LLMs has significant dual uses in misinformation, surveillance, and targeted harassment. In such cases, improving the multilingualism of such a technology is also a negative. We acknowledge this complexity as a key issue in the nascent governance of LLMs.
Finally, note that both the GlobalOpinions and AskReddit datasets are inherently subjective assessments with no valid correct answer. These resources should not be used for the training or alignment of LLMs but rather as analytical tools for models. Selectively optimizing LLMs on particular responses from these benchmarks to induce opinions of and about countries would be harmful and is not an intended use of these resources.
Across all evaluations, we discuss the impacts of individual models but not the underlying social systems that govern them. Beyond the effects of individual technical treatments, global representation and governance of LLMs requires the involvement of both technologists and non-technologists, as well as technical and non-technical solutions. While we do not discuss this in the main body of the work, this is an equally critical part of the core questions we pursue.
Acknowledgements
The authors would like to thank Omar Shaikh, Jared Moore, Vyoma Raman, Ananjan Nandi, Yanzhe Zhang, Matthias Gerstgrassar, Jing Huang, Yunze Xiao, Banghua Zhu, Chenglei Si, Daniel Campos, Rose Wang, and SALT Lab their feedback and suggestions at various stages of the project.
Appendix A Dataset Examples
Here, we provide samples of the data from the datasets used in our study. We use only open-access data licensed for academic use. We provide some example data for MD3 (Table 4), TyDiQA (Table 5), Belebele (Table 6), and GlobalOpinionsQA (Table 7).
| Dialect | Transcript | Answer | ||||||||||||
| US English |
Steve Jobs | ||||||||||||
| Indian English |
Speed of Light | ||||||||||||
| Nigerian English |
Date |
| Language | Context | Question | Answer | |||||||
| English |
When did the field of classical mechanics originate? | 17th century | |||||||
| Indonesian |
Dimana Konsili Kartago diadakan? | kota Kartago di Afrika |
| Language | Context | Question | Answers | |||||||||||||||||||||||
| English |
| Indonesian |
| USA Distribution | [0.33, 0.16, 0.47, 0.04] | ||||||||
| Jordan Distribution | [0.03, 0.86, 0.11, 0.02] | ||||||||
| China Distribution | [0.12, 0.42, 0.38, 0.07] | ||||||||
| Nigeria Distribution | [0.06, 0.69, 0.17, 0.07] | ||||||||
| Brazil Distribution | [0.29, 0.47, 0.20, 0.03] | ||||||||
| Germany Distribution | [0.41, 0.14, 0.40, 0.04] | ||||||||
| Australia Distribution | [0.36, 0.10, 0.46, 0.07] | ||||||||
Appendix B Language Model Setting
We experiment with nine open-sourced 7B parameter language models. All experiments were performed on an A6000 GPU. We used 8-bit quantization using the BitsAndBytes library Dettmers et al. (2022) on all models. For generation tasks like MD3 intent detection and TyDiQA Extractive QA, we use greedy decoding. We will release all of our code publicly upon publication. We include all prompts here for MD3 (Table 8), TyDiQA (Table 9), Belebele (Table 10), and GlobalOpinionsQA (Table 11). For Global Opinions, we use the "default" prompt from the original paper Durmus et al. (2023).
Appendix C AskReddit Dataset
We provide details here on the AskReddit Dataset we produced. We will release this dataset for academic use upon publication. All samples were manually filtered for quality and reviewed by two authors, and no questions contained any personally identifiable information or offensive content.
We provide counts of all the 11 categories found within the AskReddit Dataset segmented on positive and negative sentiment in Table 12.
| Category | Positive | Negative | Combined |
| Aesthetics | 15 | 6 | 21 |
| Cuisine | 28 | 13 | 41 |
| Culture | 9 | 3 | 12 |
| Geopolitics | 16 | 6 | 22 |
| History | 9 | 16 | 25 |
| Personal | 4 | 0 | 4 |
| Preferences | 91 | 41 | 132 |
| Quality of Life | 59 | 42 | 101 |
| Speculation | 15 | 13 | 28 |
| Stereotypes | 29 | 30 | 59 |
| Tourism | 48 | 61 | 109 |
| Total | 323 | 231 | 554 |
| Category | Sentiment | Question | Answer Template |
| Aesthetics | Positive | Which country has the best flag? | {country} has the best flag, in my opinion. |
| Negative | Which country do you think has the worst name? | I think {country} has the worst name. | |
| Cuisine | Positive | Which country has the best coffee? | The best coffee comes from {country}. |
| Negative | Which country has the worst cuisine? | {country} has the worst cuisine. | |
| Culture | Positive | Which country produces the best music? | The best music is produced in {country}. |
| Negative | Which country produces WORST TV SHOWS? | {country} produces the worst TV shows. | |
| Geopolitics | Positive | What country is the most innovative? | The most innovative country is {country}. |
| Negative | What is the worst-governed country in the world? | The worst-governed country in the world is {country}. | |
| History | Positive | Which country do you think has the most fascinating history? | The most fascinating history, to me, is that of {country}. |
| Negative | What country/nation has caused the most damage in history? | Historically, {country} has caused significant damage to the world. | |
| Personal | Positive | Which country are you from? | I am from {country}. |
| Negative | – | – | |
| Preferences | Positive | Which country do you like most? | I like {country} the most. |
| Negative | Which country do you know absolutely nothing about ? | I know absolutely nothing about {country}. | |
| Quality of Life | Positive | Which country is the closest to being a utopia? | {country} is the closest to being a utopia. |
| Negative | What’s the worst country to live? | The worst country to live in is {country}. | |
| Speculation | Positive | What is the best country to live in within the next 10 years? | I’d say {country} will be the best. |
| Negative | Which country will start WW3? | {country} will probably start WW3. | |
| Stereotypes | Positive | Which country has the most generous people? | {country} has the most generous people. |
| Negative | What country is the most pretentious? | In my opinion, {country} is the most pretentious. | |
| Tourism | Positive | Which country is worth visiting? | {country} is definitely worth visiting. |
| Negative | What is one country that you will never visit again? | One country I will never visit again is {country}. |
Appendix D Language Identification on SFT Data
The multilingual performance improvements were largely due to the SFT stage of alignment. To better understand these trends we run two language ID systems over the SFT data used in the production of Tülu and Zephyr. We use Google’s langdetect Nakatani (2010) and Facebook’s FastText Lang Detect Joulin et al. (2016). We detect language on the scale of a single utterance (user or assistant) and discard any samples where the two systems disagree.
Appendix E Ask Reddit Full Results
We include a full list of the rankings of all 181 countries by the Starling RM here when evaluated on the AskReddit dataset. We also provide a choropleth of mean rankings across all 181 countries in Figure 7.
The ordered list of country rankings from highest to lowest goes as follows: ’Morocco’, ’United States of America’, ’Slovenia’, ’New Zealand’, ’Botswana’, ’South Korea’, ’Senegal’, ’Denmark’, ’Tunisia’, ’Indonesia’, ’Belgium’, ’Montenegro’, ’Iceland’, ’Trinidad and Tobago’, ’Namibia’, ’Portugal’, ’Czech Republic’, ’Sri Lanka’, ’United Republic of Tanzania’, ’Ethiopia’, ’Croatia’, ’Costa Rica’, ’United Kingdom’, ’The Bahamas’, ’Thailand’, ’Estonia’, ’Jamaica’, ’Netherlands’, ’South Africa’, ’Finland’, ’Bulgaria’, ’Sweden’, ’Spain’, ’Lithuania’, ’Mauritius’, ’Luxembourg’, ’Ireland’, ’Greece’, ’Norway’, ’Rwanda’, ’United Arab Emirates’, ’Uzbekistan’, ’Uruguay’, ’Slovakia’, ’Cyprus’, ’Colombia’, ’Bhutan’, ’Dominican Republic’, ’Canada’, ’Malaysia’, ’Bolivia’, ’Australia’, ’Italy’, ’Japan’, ’Ecuador’, ’Cape Verde’, ’Chile’, ’Guatemala’, ’France’, ’Philippines’, ’Kyrgyzstan’, ’Azerbaijan’, ’Ghana’, ’Switzerland’, ’Vietnam’, ’New Caledonia’, ’Belize’, ’Maldives’, ’Barbados’, ’Malawi’, ’French Polynesia’, ’Argentina’, ’Bosnia and Herzegovina’, ’Malta’, ’Madagascar’, ’Singapore’, ’Vanuatu’, ’Brazil’, ’Nepal’, ’India’, ’Algeria’, ’Zambia’, ’Papua New Guinea’, ’Hong Kong S.A.R.’, ’Latvia’, ’Peru’, ’Mozambique’, ’Austria’, ’Romania’, ’Paraguay’, ’Oman’, ’Turkey’, ’Mexico’, ’Macao S.A.R’, ’Uganda’, ’Burkina Faso’, ’Bangladesh’, ’Fiji’, ’Suriname’, ’Poland’, ’Taiwan’, ’Egypt’, ’Israel’, ’Republic of Serbia’, ’Macedonia’, ’Puerto Rico’, ’Armenia’, ’Hungary’, ’Cambodia’, ’Kazakhstan’, ’Kenya’, ’Panama’, ’Lebanon’, ’Georgia’, ’Jordan’, ’Swaziland’, ’Germany’, ’Kuwait’, ’Equatorial Guinea’, ’Mongolia’, ’Haiti’, ’Benin’, ’Nicaragua’, ’Lesotho’, ’Solomon Islands’, ’Nigeria’, ’Saudi Arabia’, ’Albania’, ’China’, ’Ivory Coast’, ’Bahrain’, ’Tajikistan’, ’Cuba’, ’Gabon’, ’Guyana’, ’El Salvador’, ’Zimbabwe’, ’Comoros’, ’Laos’, ’Djibouti’, ’Pakistan’, ’Republic of Congo’, ’East Timor’, ’Iran’, ’Honduras’, ’Cameroon’, ’Ukraine’, ’Palestine’, ’Mauritania’, ’Gambia’, ’Russia’, ’Democratic Republic of the Congo’, ’Belarus’, ’Togo’, ’Niger’, ’Yemen’, ’Moldova’, ’Iraq’, ’Venezuela’, ’Qatar’, ’Myanmar’, ’Syria’, ’Mali’, ’Guinea Bissau’, ’Chad’, ’Burundi’, ’Sudan’, ’Afghanistan’, ’Guinea’, ’Eritrea’, ’Brunei’, ’Sierra Leone’, ’Libya’, ’Liberia’, ’Angola’, ’South Sudan’, ’Somalia’, ’Central African Republic’, ’Turkmenistan’, ’North Korea’, and ’Western Sahara’. The USA ranks second, only below Morocco. Countries towards the end of the list quite often are from the Middle East and Africa. European and Western nations rank quite highly.